OpenStudy (anonymous):

GIVING MEDALS!!!!! whats is the volume of the cone with a diamter of 4ft and a height of 4ft?Round to the nearest cubic foot GIVING MEDALS!!!!!

OpenStudy (anonymous):

1/3 pi r^2h 1/3 pi (2)^2(4) 1/3 pi (4)(4) 1/3 pi (16) 16pi/3

OpenStudy (anonymous):

REMINDER:YOU GET MEDALS WHEN YOU GET THE CORRECT ANSWER TO: EVERYBODY

OpenStudy (anonymous):

1/3pir^2h d=r^2 volume=1/3pi*d*h =1/3*pi*16 if mentioned pi=3.14 then volume=1/3*3.14*16 so volume=16.74666...... (or) 16.75